OpenStudy (anonymous):

Inflection points and concavity of these 2 functions: A.) 4/(1+x^2) B.) Sin(x/2) in the interval [0,4Pi]

OpenStudy (anonymous):

Inflection point by setting Y" to 0 then evaluate the x into my original function and that should be my inflection point(s) Yet cannot seem to get it right.

OpenStudy (anonymous):

F'(x) of 4/(1+x^2) = -8x/(1+x^2)^2 Correct?

OpenStudy (anonymous):

The F" of 4/(1+x^2) = -8+24x^2/(1+x^2)3 Correct?

OpenStudy (anonymous):

So the result of that set to 0 is equal to -1/SQROOT(3) and 1/SQROOT(3) Correct???

OpenStudy (anonymous):

\[\frac{ 4 }{ 1+x ^{2} } = \frac{ 4 }{ 1+(\frac{ 1 }{ \sqrt{3} }) } \]

OpenStudy (anonymous):

the \[(\frac{ 1 }{ \sqrt{3} })^{2}\] Should have looked like that

OpenStudy (anonymous):

Should equal positive 3 in both cases so is my Inflection point: \[(\frac{ 1 }{\sqrt{3} } , 3) \] and \[(-\frac{ 1 }{ \sqrt{3} }, 3)\] or is it just one of those?

OpenStudy (schrodingers_cat):

It is both does the answer come out wrong?

OpenStudy (anonymous):

Not sure it was a hand giving problem to study with, wondering for my own sake if it is correct or not and out of 161 people here only 1 response. Thank you

OpenStudy (schrodingers_cat):